# Eigenvalues of Heckman-Polychronakos operators

### v1

@@ 5. Sums of eigenvalues
In general, the eigenvalue of the operator $\mathcal{P}_{m}$ on the function $F_{\lambda,\tau,i}$ can be expressed as a root of a polynomial equation (of degree at most $\dim(\tau)$) with coefficients being polynomials in $\lambda_{i}$ and $\theta$; therefore, this eigenvalue can be quite complicated, see, e.g. (33) in Section 6. We found that the sums of all eigenvalues corresponding to the same $\lambda$ and $\tau$, are much more manageable.
@@ 5.1. Distinct degrees of variables
For an irreducible representation of $S_{N}$ labeled by $\tau\vdash N$, we let $\dim\tau$ be its dimension and let $\chi^{\tau}(g)$ be its character. In particular, $\dim(\tau)=\chi^{\tau}(\mathrm{Id})$. We also recall the notation $h_{m}^{(r)}$ from (3).
@@
(13) $\sum_{i=1}^{\dim V_{\lambda,\tau}}\mathrm{eig}_{m}(\lambda,\tau,i)=\dim\tau\sum_{k=1}^{\min(m+1,N)}(-\theta)^{k-1}h^{(k)}_{m+1-k}\bigl{(}\ell_{1},\dots,\ell_{N}\bigr{)}\chi^{\tau}\bigl{(}(1,2,\dots,k)),$
where $(1,2,\dots,k)$ is the permutation having a single $k$-cycle and $N-k$ fixed points.
When $\tau=(N)$, $\dim V_{\lambda,\tau}=1$ and the expression (13) turns into (5); this eigenvalue corresponds to the symmetric eigenfunction. When $\tau=1^{N}$, also $\dim V_{\lambda,\tau}=1$ and the expression (13) turns into a formula similar to (5) but with $(-\theta)^{k-1}$ replaced with $\theta^{k-1}$; this eigenvalue corresponds to a skew-symmetric eigenfunction.
@@
For the proof of Theorem 5.1 we use $T_{i}$ of (9). We also need an additional object from representation theory of $S_{N}$. For each irreducible representation $\tau$ of $S_{N}$, denote
@@
[Ser77, Theorem 8] says that in each representation of $S_{N}$, $\pi_{\tau}$ acts as an orthogonal projector (with respect to an $S_{N}$–invariant scalar product) onto isotypical component of $\tau$.
@@
Recall from the proof of Theorem 4.3 that $\mathrm{eig}_{m}(\lambda,\tau,i)$ are eigenvalues of $(T_{1}^{m}+\dots+T_{n}^{m})$, corresponding to eigenfunctions in $V_{\lambda;\tau}$. Hence, $\mathrm{eig}_{m}(\lambda,\tau,i)$ are eigenvalues of the restriction of $(T_{1}^{m}+\dots+T_{n}^{m})$ onto $V_{\lambda;\tau}$, which is the same as eigenvalues of $\pi_{\tau}(T_{1}^{m}+\dots+T_{n}^{m})\pi_{\tau}$. The sum of the eigenvalues coincides with trace, and we get $\mathrm{Trace}_{V_{\lambda}}\left[\pi_{\tau}(T_{1}^{m}+\dots+T_{n}^{m})\pi_{\tau}\right]$. Since $\mathrm{Trace}[AB]=\mathrm{Trace}[BA]$, we can move one $\pi_{\tau}$ inside the trace from right to the left, and then using the projecting property $\pi_{\tau}\pi_{\tau}=\pi_{\tau}$, we arrive at (16). ∎
@@
We use Lemma 5.2 and compute $\mathrm{Trace}_{V_{\lambda}}\left[\pi_{\tau}T_{1}^{m}\right]$ combining (9) with (15). The computation is based on the observation that $V_{\lambda}$ is isomorphic to the regular representation of $S_{N}$, and therefore in the trace computation all products of $g$ and $(1,j)$ not equal to the identity permutation do not give any contribution.
@@
where $j_{1},\dots,j_{k-1}$ are distinct indices, such that each of them corresponds to distinct $\gamma_{j_{k-1}}>\gamma_{j_{k-2}}>\dots>\gamma_{j_{1}}>\gamma_{1}$. The transpositions $(1,j_{a})$ correspond to the second term in (9) and the ordering appears because the operator $T_{1}$ can only increase the degree of $x_{1}$, but never decrease. On the other hand, the prefactor $p_{k;q}$ collects all the factors coming from the first term; it is a polynomial in various $\theta$-shifted $\gamma_{i}$ of degree $(m+1-k)$, which depends on the collection $\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}$ and on $\theta$. Note also that the product of $k-1$ transpositions becomes a $k$–cycle
@@
and note that when we plug into $\mathrm{Trace}_{V_{\lambda}}\left[\pi_{\tau}T_{1}^{m}\right]$, and multiply the sums, only the terms with $\tilde{g}=\mathrm{Id}$ give non-zero contributions. We also note that the trace is a central function and its values on all $k$–cycles are the same. Therefore,
@@
Next, we notice that the first term in (9) is always equal to of $\ell_{a}$ and therefore matches the diagonal element of the matrix in (19). Similarly, the $-\theta(i,j)$ terms in (9) match the off-diagonal $(-\theta)$ elements in the matrix in (19). Hence, we see that the combinatorics of rising the matrix to the $m$–th power in (19) and raising $T_{i}$ to the $m$–th power is exactly the same. Hence, the identity (19).
It remains to identify the right-hand side of (19) with $h^{(k)}_{m+1-k}\bigl{(}\ell_{1},\dots,\ell_{N}\bigr{)}$, which is immediate from the definition (3): the $k$–element subset $\{i_{1},\dots,i_{k}\}$ encodes those diagonal elements of the matrix in (19), which enter into the computation of the $m$–th power.
@@ 5.2. General degrees
Given an integer $0\leq p\leq N$, a multiplicity composition (i.e. a sequence of $p$ positive integers) $\mathbf{n}=(n_{1},n_{2},\dots,n_{p}>0)$ with $\sum_{i=1}^{p}n_{i}=N$, and a degree sequence $\mathbf{d}=(d_{1}>d_{2}>\dots>d_{p}\geq 0)$, we introduce a partition
@@
For any $k$–element subset $\mathcal{A}\subset\{1,\dots,p\}$ of the form $\{a_{1}<a_{2}<\dots<a_{k}\}$, we introduce averaged characters also called “spherical functions”:
@@
~$\tilde{\ell}_{i}=d_{i}+\theta(n_{i+1}+n_{i+2}+\dots+n_{p}),\qquad 1\leq i\leq p.$
Note that {~$\{\tilde{\ell}_{i}\}_{i=1}^{p}$ is a subset of $\{\ell_{i}=\lambda_{i}+\theta(N-i)\}_{i=1}^{N}$. Finally, we define complete homogeneous functions in subsets of these numbers:
@@
Note that for $p=N$, the function $h^{(k)}_{m}$, as in (3), is the sum of $h^{\mathcal{A}}$ over all $k$–element subsets.
@@
If $\tau=(N)$ is the trivial representation, then the eigenfunction is the Jack polynomials (see Lemma 7.4 for some details) and simultaneously $\chi^{\tau}[\mathcal{A};\mathbf{n}]=1$. Yet, the expression we get in (23) is slightly different from the one of Theorem 3.1, and we prove their equivalence in Lemma 5.5 below.
If $\tau=(1^{N})$ is the sign representation, then $\chi^{\tau}[\mathcal{A};\mathbf{n}]=0$, unless $\lambda$ has all distinct parts. In fact, there is no $(1^{N})$–type component in $V_{\lambda}$ for $\lambda$ with non-trivial multiplicities of parts.
@@
terms. Absorbing the factorials from the denominator into the character summation (they are present in (21)) and cancelling $N!$ with $(N-1)!$ in denominator in (25), we are left with a sum involving a single fixed $\gamma$ and $\pi_{\tau}\sum_{i=1}^{N}(T_{i})^{m}$.
@@
and the corresponding degrees $\gamma_{j_{k-1}},\gamma_{j_{k-2}},\dots,\gamma_{1}$ are arranged in the decreasing order, which correspond to increasing indices of $d_{i}$ — matching the cycle $\mathfrak{c}$ used in (21). The stabilizer of $\gamma$, i.e. the subgroup $\tilde{g}\in S_{N}\mid\tilde{g}\gamma=\gamma$ is not necessary $S_{n_{1}}\times S_{n_{2}}\times\dots S_{n_{p}}$, but it is isomorphic to this group, with isomorphism obtained by renaming $\{1,2,\dots,N\}$. Since the characters are central, this leads to the same function as (21).
It remains to identify the sums of $p_{k;m}[\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}]$. Note that when we act with $(T_{i})^{m}$ using its definition (9), the factors $\left(\gamma_{i}+\theta\#\{j\mid\gamma_{j}<\gamma_{i}\}\right)$ correspond to various ~$\tilde{\ell}_{a}$. On the other hand, each application of a term from $\theta\sum_{j\mid\gamma_{j}>\gamma_{i}}(i,j)x^{\gamma}$, leads to addition of an additional index ${j_{a}}$. In this index, only the value of $\gamma_{j_{a}}$ matters. The set $\mathcal{A}$ in (23) is identified with $\{\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}\}$ in (25), leading to the combinatorial factor $\prod_{a\in\mathcal{A}}n_{a}$, counting the number of ways to choose indices $\gamma_{j_{a}}$ corresponding to $\mathcal{A}$ (the $n_{a}$ corresponding to $\gamma_{1}$ should be also accounted, as it arises from the first application of one of the operators $T_{i}$ in $\sum_{i=1}^{N}(T_{i})^{m}$). Once the set $\mathcal{A}$ is fixed, the sums of products of ~$\tilde{\ell}_{a}$ coming from $\left(\gamma_{i}+\theta\#\{j\mid\gamma_{j}<\gamma_{i}\}\right)$ in $T_{i}$ combine precisely into $h^{\mathcal{A}}_{m+1-k}$ of (23). ∎
Lemma 5.5.
@@
The formula (5) can be rewritten in terms of the $m$–th power of $N\times N$ matrix as
@@
where $\ell_{i}=\lambda_{i}+\theta(N-i)$. On the other hand, the formula (23) can rewritten in terms of $m$–th power of $p\times p$ matrix as
@@
where ~$\tilde{\ell}_{i}$ are a subset of $\ell_{i}$ given by ~$\tilde{\ell}_{i}=d_{i}+\theta(n_{i+1}+n_{i+2}+\dots+n_{p})$ in the notation (20). Our task is to prove (26)=(27). One can do this directly, but our approach is to proceed through generating functions. On the (26) side, the generating function was computed in Proposition 3.2. Let $\widetilde{\mathrm{eig}}_{m}(\lambda)$ denote the eigenvalues computed by (27). Repeating the argument of Proposition 3.2, we have
@@ 5.3. $(N-1,1)$ isotype
The irreducible representation of $S_{N}$ of type $\tau=(N-1,1)$ is the standard $(N-1)$–dimensional representation in the space of vectors $(x_{1},\dots,x_{N})$ with $\sum_{i=1}^{N}x_{i}=0$ and action $\sigma\circ(x_{1},\dots,x_{N})=(x_{\sigma^{-1}(1)},\dots,x_{\sigma^{-1}(N)})$. Evaluating the character as trace, we get an expression in terms of the number of the fixed points:
@@
where $\mathfrak{c}$ is a $|\mathcal{A}|$-cycle permuting the groups correspinding to the elements $a\in\mathcal{A}$. Changing the order of summation in (29), we get
(30) $\sum_{q=1}^{p}\,\sum_{i=n_{1}+\dots+n_{q-1}+1}^{n_{1}+\dots+n_{q}}\,\sum_{\tilde{g}\in S_{n_{1}}\times S_{n_{2}}\times\dots\times S_{n_{p}}}\mathbf{1}_{\tilde{g}\cdot\mathfrak{c}(i)=i}-\prod_{j=1}^{p}(n_{j})!\\ =\sum_{q=1}^{p}\left[\mathbf{1}_{q\in\mathcal{A}}(n_{q}-1)+\mathbf{1}_{q\not\in\mathcal{A}}n_{q}\right](n_{q}-1)!\left[\prod_{j\neq q}(n_{j})!\right]-\prod_{j=1}^{p}(n_{j})!,\\ =(p-1)\prod_{j=1}^{p}(n_{j})!-\sum_{q=1}^{p}\mathbf{1}_{q\in\mathcal{A}}(n_{q}-1)!\left[\prod_{j\neq q}(n_{j})!\right].$

### v2

@@ 5. Sums of eigenvalues
In general, the eigenvalue of the operator $\mathcal{P}_{m}$ on the polynomial $F_{\lambda,\tau,i}$ can be expressed as a root of a polynomial equation (of degree at most $\dim(\tau)$) with coefficients being polynomials in $\lambda_{i}$ and $\theta$; therefore, this eigenvalue can be quite complicated, see, e.g. (32) in Section 6. We found that the sums of all eigenvalues corresponding to the same $\lambda$ and $\tau$, are much more manageable.
@@ 5.1. Distinct degrees of variables
For an irreducible representation of $S_{N}$ labeled by $\tau\vdash N$, we let $\dim\tau$ be its dimension and let $\chi^{\tau}(g)$ be its character. In particular, $\dim(\tau)=\chi^{\tau}(\mathrm{Id})$. We recall the notation $h_{m}^{(r)}$ from (3). We also recall that $(1,2,\dots,k)$ is the permutation having a single $k$-cycle and NkN-k fixed points $\{k+1,\dots,N\}$.
@@
(13) $\sum_{i=1}^{\dim V_{\lambda,\tau}}\mathrm{eig}_{m}(\lambda,\tau,i)=\dim\tau\sum_{k=1}^{\min(m+1,N)}(-\theta)^{k-1}h^{(k)}_{m+1-k}\bigl{(}\ell_{1},\dots,\ell_{N}\bigr{)}\chi^{\tau}\bigl{(}(1,2,\dots,k)).$
When $\tau=(N)$, $\dim V_{\lambda,\tau}=1$ and the expression (13) turns into (5); this eigenvalue corresponds to the symmetric eigenfunction. When $\tau=(1^{N})$, also $\dim V_{\lambda,\tau}=1$ and the expression (13) turns into a formula similar to (5) but with $(-\theta)^{k-1}$ replaced with $\theta^{k-1}$; this eigenvalue corresponds to a skew-symmetric eigenfunction.
@@
For the proof of Theorem 5.1 we use $T_{i}$ of (9). We also need an additional object from the representation theory of $S_{N}$. For each irreducible representation $\tau$ of $S_{N}$, denote
@@
[Ser77, Theorem 8] says that in each representation of $S_{N}$, $\pi_{\tau}$ acts as an orthogonal projector (with respect to an $S_{N}$-invariant scalar product) onto the isotypical component of $\tau$.
@@
Recall from the proof of Theorem 4.3 that $\mathrm{eig}_{m}(\lambda,\tau,i)$ are eigenvalues of $(T_{1}^{m}+\dots+T_{n}^{m})$, corresponding to eigenfunctions in $V_{\lambda;\tau}$. Hence, $\mathrm{eig}_{m}(\lambda,\tau,i)$ are eigenvalues of the restriction of $(T_{1}^{m}+\dots+T_{n}^{m})$ onto $V_{\lambda;\tau}$, which is the same as eigenvalues of $\pi_{\tau}(T_{1}^{m}+\dots+T_{n}^{m})\pi_{\tau}$. The sum of the eigenvalues coincides with the trace of the operator, and we get $\mathrm{Trace}_{V_{\lambda}}\left[\pi_{\tau}(T_{1}^{m}+\dots+T_{n}^{m})\pi_{\tau}\right]$. Since $\mathrm{Trace}[AB]=\mathrm{Trace}[BA]$, we can move one $\pi_{\tau}$ inside the trace from right to the left, and then using the projecting property $\pi_{\tau}\pi_{\tau}=\pi_{\tau}$, we arrive at (16). ∎
@@
We use Lemma 5.2 and compute $\mathrm{Trace}_{V_{\lambda}}\left[\pi_{\tau}T_{1}^{m}\right]$ combining (9) with (15). The computation is based on the observation that $V_{\lambda}$ is isomorphic to the regular representation of $S_{N}$, because for distinct permutations $\sigma\in S_{N}$, their actions on $x_{1}^{\lambda_{1}}\cdots x_{N}^{\lambda_{N}}$ give distinct monomials. Therefore, in the trace computation all products of $g$ and $(1,j)$ not equal to the identity permutation do not give any contribution.
@@
where $j_{1},\dots,j_{k-1}$ are distinct indices, such that each of them corresponds to distinct $\gamma_{j_{k-1}}>\gamma_{j_{k-2}}>\dots>\gamma_{j_{1}}>\gamma_{1}$. The transpositions $(1,j_{a})$ correspond to the second term in (9) and the ordering appears because the operator $T_{1}$ can only increase the degree of $x_{1}$, but never decrease. On the other hand, the prefactor $p_{k;q}$ collects all the factors coming from the first term; it is a polynomial in various $\theta$-shifted $\gamma_{i}$ of degree (m+1k)(m+1-k), which depends on the collection $\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}$ and on $\theta$. Note also that the product of k1k-1 transpositions becomes a $k$-cycle
@@
and note that when we plug into $\mathrm{Trace}_{V_{\lambda}}\left[\pi_{\tau}T_{1}^{m}\right]$, and multiply the sums, only the terms with $\tilde{g}=\mathrm{Id}$ give non-zero contributions. We also note that the trace is a central function and its values on all $k$-cycles are the same. Therefore,
@@
Next, we notice that the first term in (9) is always equal to $\ell_{a}$ and therefore matches the diagonal element of the matrix in (19). Similarly, the $\theta(i,j)$ terms in (9) match the off-diagonal $(-\theta)$ elements in the matrix in (19). Hence, we see that the combinatorics of rising the matrix to the $m$-th power in (19) and raising $T_{i}$ to the $m$-th power is exactly the same. Hence, the identity (19).
It remains to identify the right-hand side of (19) with $h^{(k)}_{m+1-k}\bigl{(}\ell_{1},\dots,\ell_{N}\bigr{)}$, which is immediate from the definition (3): the $k$-element subset $\{i_{1},\dots,i_{k}\}$ encodes those diagonal elements of the matrix in (19), which enter into the computation of the $m$-th power.
@@ 5.2. General degrees
Given an integer $1\leq p\leq N$, a multiplicity composition (i.e. a sequence of $p$ positive integers) $\mathbf{n}=(n_{1},n_{2},\dots,n_{p}>0)$ with $\sum_{i=1}^{p}n_{i}=N$, and a degree sequence $\mathbf{d}=(d_{1}>d_{2}>\dots>d_{p}\geq 0)$, we introduce a partition
@@
For any $k$-element subset $\mathcal{A}\subset\{1,\dots,p\}$ of the form $\{a_{1}<a_{2}<\dots<a_{k}\}$, we introduce averaged characters also called “spherical functions”:
@@
$\tilde{\ell}_{i}=d_{i}+\theta(n_{i+1}+n_{i+2}+\dots+n_{p}),\qquad 1\leq i\leq p.$
Note that $\{\tilde{\ell}_{i}\}_{i=1}^{p}$ is a subset of $\{\ell_{i}=\lambda_{i}+\theta(N-i)\}_{i=1}^{N}$. Finally, we define complete homogeneous polynomials in subsets of these numbers:
@@
Note that for $p=N$, the polynomial $h^{(k)}_{m}$, as in (3), is the sum of $h^{\mathcal{A}}$ over all $k$-element subsets.
@@
If $\tau=(N)$ is the trivial representation, then the eigenfunction is the Jack polynomial (see Lemma 7.4 for some details) and simultaneously $\chi^{\tau}[\mathcal{A};\mathbf{n}]=1$. Yet, the expression we get in (23) is slightly different from the one of Theorem 3.1, and we prove their equivalence in Lemma 5.5 below.
If $\tau=(1^{N})$ is the sign representation, then $\chi^{\tau}[\mathcal{A};\mathbf{n}]=0$, unless $\lambda$ has all distinct parts. In fact, there is no $(1^{N})$-type component in $V_{\lambda}$ for $\lambda$ with non-trivial multiplicities of parts.
@@
terms. Absorbing the factorials from the denominator into the character summation (they are present in (21)) and cancelling $N!$ with (N1)!(N-1)! in denominator in (25), we are left with a sum involving a single fixed $\gamma$ and $\pi_{\tau}\sum_{i=1}^{N}(T_{i})^{m}$.
@@
and the corresponding degrees $\gamma_{j_{k-1}},\gamma_{j_{k-2}},\dots,\gamma_{1}$ are arranged in the decreasing order, which correspond to increasing indices of $d_{i}$ — matching the cycle $\mathfrak{c}$ used in (21). The stabilizer of $\gamma$, i.e. the subgroup $\{\tilde{g}\in S_{N}\mid\tilde{g}\gamma=\gamma\}$ is not necessarily $S_{n_{1}}\times S_{n_{2}}\times\dots S_{n_{p}}$, but it is isomorphic to this group, with isomorphism obtained by renaming $\{1,2,\dots,N\}$. Since the characters are central, this leads to the same function as (21).
It remains to identify the sums of $p_{k;m}[\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}]$. Note that when we act with $(T_{i})^{m}$ using its definition (9), the factors $\left(\gamma_{i}+\theta\#\{j\mid\gamma_{j}<\gamma_{i}\}\right)$ correspond to various $\tilde{\ell}_{a}$. On the other hand, each application of a term from $\theta\sum_{j\mid\gamma_{j}>\gamma_{i}}(i,j)x^{\gamma}$, leads to addition of an additional index $j_{a}$}. In this index, only the value of $\gamma_{j_{a}}$ matters. The set $\mathcal{A}$ in (23) is identified with $\{\gamma_{1},\gamma_{j_{1}},\dots,\gamma_{j_{k-1}}\}$ in (25), leading to the combinatorial factor $\prod_{a\in\mathcal{A}}n_{a}$, counting the number of ways to choose indices $\gamma_{j_{a}}$ corresponding to $\mathcal{A}$ (the $n_{a}$ corresponding to $\gamma_{1}$ should be also accounted, as it arises from the first application of one of the operators $T_{i}$ in $\sum_{i=1}^{N}(T_{i})^{m}$). Once the set $\mathcal{A}$ is fixed, the sums of products of $\tilde{\ell}_{a}$ coming from $\left(\gamma_{i}+\theta\#\{j\mid\gamma_{j}<\gamma_{i}\}\right)$ in $T_{i}$ combine precisely into $h^{\mathcal{A}}_{m+1-k}$ of (23). ∎
Lemma $5.$
@@
The formula (5) can be rewritten in terms of the $m$-th power of $N\times N$ matrix as
@@
where $\ell_{i}=\lambda_{i}+\theta(N-i)$. On the other hand, the formula (23) can rewritten in terms of $m$-th power of $p\times p$ matrix as
@@
where $\tilde{\ell}_{i}$ are a subset of $\ell_{i}$ given by $\tilde{\ell}_{i}=d_{i}+\theta(n_{i+1}+n_{i+2}+\dots+n_{p})$ in the notation (20). Our task is to prove (26)=(27). One can do this directly, but our approach is to proceed through generating functions. On the (26) side, the generating function was computed in Proposition 3.2. Let $\widetilde{\mathrm{eig}}_{m}(\lambda)$ denote the eigenvalues computed by (27). Repeating the argument of Proposition 3.2, we have
@@
The irreducible representation of $S_{N}$ of type $\tau=(N-1,1)$ is the standard (reflection) (N1)(N-1)-dimensional representation in the space of vectors $(x_{1},\dots,x_{N})$ with $\sum_{i=1}^{N}x_{i}=0$ and action $\sigma\circ(x_{1},\dots,x_{N})=(x_{\sigma^{-1}(1)},\dots,x_{\sigma^{-1}(N)})$. Evaluating the character as trace, we get an expression in terms of the number of the fixed points:
@@
where $\mathfrak{c}$ is a $|\mathcal{A}|$-cycle permuting the groups corresponding to the elements $a\in\mathcal{A}$. Changing the order of summation in (29), we get
$\sum_{q=1}^{p}\,\sum_{i=n_{1}+\dots+n_{q-1}+1}^{n_{1}+\dots+n_{q}}\,\sum_{\tilde{g}\in S_{n_{1}}\times S_{n_{2}}\times\dots\times S_{n_{p}}}\mathbf{1}_{\tilde{g}\cdot\mathfrak{c}(i)=i}-\prod_{j=1}^{p}(n_{j})!.$
In the internal sum over $\tilde{g}$, for the symmetric groups $S_{n_{j}}$ which permute the elements other than $\mathfrak{c}(i)$ there are no restrictions and all $(n_{j})!$ permutations are possible. However, if $\mathfrak{c}(i)$ belongs to the indices permuted by $S_{n_{j}}$, then there are two options: if $\mathfrak{c}(i)=i$, then $(n_{j}-1)!$ permutations satisfy the restriction $\tilde{g}\cdot\mathfrak{c}(i)=i$, otherwise no permutations satisfy the restriction. Hence, summing over $\tilde{g}$ and then $i$, we get
$\sum_{q=1}^{p}\left[\mathbf{1}_{q\in\mathcal{A}}(n_{q}-1)+\mathbf{1}_{q\not\in\mathcal{A}}n_{q}\right](n_{q}-1)!\left[\prod_{j\neq q}(n_{j})!\right]-\prod_{j=1}^{p}(n_{j})!.$
The last expression can be transformed into
$(p-1)\prod_{j=1}^{p}(n_{j})!-\sum_{q=1}^{p}\mathbf{1}_{q\in\mathcal{A}}(n_{q}-1)!\left[\prod_{j\neq q}(n_{j})!\right].$